Assume your organization has the following inventory changes during the year:

Beginning Inventory 15 units Valued at $10,000 each

February purchases 13 units at $11,500 each

June purchases 20 units at $12,000 each

Total units used 42

Calculate the value of the ending inventory and the value of the inventory used (the inventory expense) for the year, using both the FIFO and the LIFO method of cost-flow.
Answer #1

FIFO

Total units Available for sale = 15 + 13 + 20 = 48 units

Total units used = 42 units

Cost of Ending Inventory = (48 - 42) * $12,000 = $72,000

Cost of Ending Inventory = $72,000

Value of Inventory used = (15 * $10,000) + ($13 * $11,500) + (14 * $12,000)

Value of Inventory used = $467,500

LIFO

Total units Available for sale = 15 + 13 + 20 = 48 units

Total units used = 42 units

Cost of Ending Inventory = (48 - 42) * $10,000 = $60,000

Cost of Ending Inventory = $60,000

Value of Inventory used = (20 * $12,000) + ($13 * $11,500) + (9 * $10,000)

Value of Inventory used = $479,500